OpenStudy (anonymous):

HELPP!!!! The exterior angles of a quadrilateral measure 3x + 10, 2x - 60, 3x + 23, and 2x - 13. Find the measure of each angle.?

OpenStudy (anonymous):

A) -20°, 87°, 145°, and 243° Eliminate Reactivate B) -16°, 31°, 76°, and 89° Eliminate Reactivate C) 23°, 25°, 38°, and 50° Eliminate Reactivate D) 20°, 67°, 130°, and 143°

Directrix (directrix):

What is the sum of the exterior angles of a polygon?

Directrix (directrix):

3x + 10 + 2x - 60 + 3x +23 + 2x -13 = 360

Directrix (directrix):

Are you there?

OpenStudy (anonymous):

YEAHH THATS RIGHT

Directrix (directrix):

We have to solve this equation: What's this eliminate - reactivate stuff?

Directrix (directrix):

10x - 40 = 360 x = 40 3x + 10= 3(40) + 10 = 2x - 60= 2(40) - 60- 3x + 23= 3(40) + 23= 2x - 13 =2(40) - 30 =

Directrix (directrix):

Anybody here?